# Forget Vectors at Play: Universal Input Perturbations Driving Machine Unlearning in Image Classification

### v2

@@ Appendix A Generalization of MU to Forget Data Shifts
Appendix B Parameter Sensitivity Analysis: Prediction Margin $\tau$ in Forget Vector Loss
@@
Appendix C Ablation Studies
C.1 Component Analysis: $\lambda_{1}$ and $\lambda_{2}$
@@
To verify the impact of each key component in the optimization problem (4) of the main paper, we analyze the nonnegative trade-off parameters $\lambda_{1}$ and $\lambda_{2}$ in Fig. A3. As can be seen in Fig. A3-(a) and (c), setting the retain loss regularization parameter $\lambda_{1}$ to $3$ for CIFAR-10 and $5$ for ImageNet-10 in class-wise forgetting, along with the $\ell_{2}$-norm regularization parameter $\lambda_{2}=1$, enables our proposed method to achieve the highest average performance. Meanwhile, for the random data forgetting scenario, setting both $\lambda_{1}$ and $\lambda_{2}$ to $1$ yields the best average performance for CIFAR-10 and ImageNet-10.
C.2 Run-time Efficiency
@@
Table A1: The computational efficiency overview of various MU methods for image classification under two unlearning scenarios on CIFAR-10 and ImageNet-10 using ResNet-18 and VGG-16, respectively. Results are reported in terms of RTE (run-time efficiency) measured in minutes for the overall training phase.
MU Method CIFAR-10 ImageNet-10
Class-wise Forgetting Random Data Forgetting Class-wise Forgetting Random Data Forgetting
Retrain $81.30$ $103.71$ $118.89$ $116.10$
FT (Warnecke et al., 2021) $4.56$ $2.28$ $2.99$ $2.95$
RL (Golatkar et al., 2020) $5.61$ $4.75$ $3.58$ $3.59$
GA (Thudi et al., 2022a) $1.16$ $1.15$ $0.30$ $0.31$
NegGrad+ (Kurmanji et al., 2024a) $4.72$ $4.22$ $3.21$ $3.17$
SalUn (Fan et al., 2023) $2.81$ $3.01$ $2.56$ $2.85$
SCRUB (Kurmanji et al., 2024b) $1.23$ $1.53$ $1.62$ $1.55$
Ours $9.48$ $10.75$ $24.60$ $25.99$
To evaluate the computational efficiency in MU, we compare the runtime costs of forget vector calculation with those of other model-based unlearning methods. Following (Fan et al., 2023), we adopt the evaluation metric run-time efficiency (RTE), i.e., the computation time of applying an MU method in minutes. Notably, all evaluations are conducted on the same computational environment utilizing an NVIDIA A6000 GPU, using a consistent batch size across all MU methods to maintain fair and reliable comparisons. The corresponding results can be found in Tab. A1. By comparing the training time of model-based methods with our proposed approach, we observe that the forget vector method requires higher computational time than model-based unlearning baselines but remains faster than retraining. This increased time is attributed to the larger number of iterations needed to identify the desired universal input perturbations. However, it is worth noting that the forget vector method optimizes fewer parameters, owing to its input-level design, compared to model-based methods. This trade-off highlights the efficiency of the forget vector in terms of parameter scale despite its higher iteration count.
Appendix D Additional Visualization through An Input Saliency Lens

### v4

@@ Appendix B Generalization of MU to Forget Data Shifts
Appendix C Parameter Sensitivity Analysis: Prediction Margin $\tau$ in Forget Vector Loss
@@
Appendix D Performance on ImageNet-10 using VGG-16
[…]
Class-wise Forgetting, ImageNet-10, VGG-16 Random Data Forgetting, ImageNet-10, VGG-16
Retrain $100.00_{\pm{0.00}}$($0.00$) $100.00_{\pm{0.00}}$($0.00$) $99.66_{\pm{0.16}}$($0.00$) $97.11_{\pm{0.82}}$($0.00$) $\mathbf{0.00}$ $4.05_{\pm{0.45}}$($0.00$) $6.60_{\pm{1.07}}$($0.00$) $99.48_{\pm{0.07}}$($0.00$) $96.33_{\pm{0.38}}$($0.00$) $\mathbf{0.00}$
FT $39.66_{\pm{4.73}}$($60.34$) $55.76_{\pm{7.26}}$($44.24$) $99.78_{\pm{0.03}}$($0.13$) $97.27_{\pm{0.35}}$($2.35$) $\mathbf{26.77}$ $1.35_{\pm{0.32}}$($2.70$) $4.67_{\pm{1.61}}$($1.93$) $99.36_{\pm{0.28}}$($0.12$) $96.54_{\pm{0.59}}$($0.21$) $\mathbf{1.24}$
RL $76.58_{\pm{11.64}}$($23.42$) $46.04_{\pm{33.71}}$($53.96$) $99.28_{\pm{0.20}}$($0.63$) $96.91_{\pm{0.55}}$($1.99$) $\mathbf{20.00}$ $2.96_{\pm{0.42}}$($1.09$) $12.85_{\pm{4.25}}$($6.25$) $99.19_{\pm{0.16}}$($0.29$) $95.50_{\pm{0.90}}$($0.83$) $\mathbf{2.12}$
GA $46.61_{\pm{6.11}}$($53.39$) $49.15_{\pm{9.36}}$($50.85$) $99.35_{\pm{0.11}}$($0.56$) $95.60_{\pm{0.22}}$($0.68$) $\mathbf{26.37}$ $0.18_{\pm{0.04}}$($3.87$) $2.97_{\pm{1.51}}$($3.63$) $99.86_{\pm{0.01}}$($0.38$) $97.47_{\pm{0.09}}$($1.14$) $\mathbf{2.23}$
NegGrad+ $49.56_{\pm{34.87}}$($50.44$) $64.27_{\pm{27.33}}$($35.73$) $99.08_{\pm{1.18}}$($0.58$) $96.47_{\pm{1.25}}$($0.64$) $\mathbf{21.84}$ $0.60_{\pm{0.44}}$($3.45$) $3.90_{\pm{2.22}}$($2.70$) $99.68_{\pm{0.28}}$($0.20$) $97.12_{\pm{0.45}}$($0.79$) $\mathbf{1.79}$
SalUn $95.13_{\pm{1.79}}$($4.87$) $97.24_{\pm{0.17}}$($2.76$) $96.33_{\pm{0.25}}$($3.33$) $96.18_{\pm{1.10}}$($0.93$) $\mathbf{2.97}$ $1.15_{\pm{0.40}}$($2.90$) $3.56_{\pm{1.12}}$($3.04$) $98.97_{\pm{1.56}}$($0.51$) $95.42_{\pm{2.10}}$($0.91$) $\mathbf{1.84}$
SCRUB $99.12_{\pm{0.14}}$($0.88$) $98.01_{\pm{0.51}}$($1.99$) $99.75_{\pm{0.03}}$($0.09$) $97.24_{\pm{0.16}}$($0.13$) $\mathbf{0.77}$ $0.18_{\pm{0.08}}$($3.87$) $2.80_{\pm{1.33}}$($3.80$) $99.88_{\pm{0.03}}$($0.40$) $97.36_{\pm{0.15}}$($1.03$) $\mathbf{2.28}$
Class-F $71.19_{\pm{3.50}}$($28.81$) $63.57_{\pm{4.07}}$($36.43$) $61.55_{\pm{4.25}}$($38.11$) $59.40_{\pm{4.59}}$($37.71$) $\mathbf{35.27}$ n/a n/a n/a n/a n/a
SSD $91.50_{\pm{0.96}}$($8.50$) $91.30_{\pm{8.70}}$($8.70$) $99.48_{\pm{0.18}}$($0.18$) $97.00_{\pm{0.11}}$($0.11$) $\mathbf{4.38}$ $0.88_{\pm{0.04}}$($3.17$) $5.35_{\pm{0.45}}$($1.25$) $99.36_{\pm{0.15}}$($0.12$) $96.40_{\pm{0.40}}$($0.07$) $\mathbf{1.15}$
Ours $87.23_{\pm{6.55}}$($12.77$) $91.41_{\pm{5.90}}$($8.59$) $94.77_{\pm{1.16}}$($5.14$) $94.04_{\pm{1.29}}$($0.88$) $\mathbf{6.85}$ $2.27_{\pm{0.50}}$($1.78$) $6.13_{\pm{1.40}}$($0.47$) $98.29_{\pm{0.32}}$($1.19$) $95.82_{\pm{0.48}}$($0.51$) $\mathbf{0.99}$
Table A1: Performance overview of various MU methods for image classification under two unlearning scenarios on ImageNet-10 using VGG-16. Since Class-F (Kodge et al., 2024) is specifically designed for class-wise forgetting, its results do not apply to random data forgetting scenarios (n/a). Other results are reported in the format $a_{\pm b}$, where $a$ is the mean and $b$ denotes standard deviation $b$ over $10$ independent trials. The performance gap against Retrain is indicated in ($\bullet$), where a lower value is better. $\uparrow$ (or $\downarrow$) indicates that a higher (or lower) value is better. The best performance for each metric is highlighted in green , while the second-best performance is highlighted in red .
In table A1, we provide additional performance overview of different MU methods under two unlearning scenarios on ImageNet-10 using VGG-16. Considering the gain in unlearning effectiveness, the Avg. Gap with Retrain shows that the forget vector remains competitive, except for class-wise forgetting on (ImageNet-10, VGG-16). In this case, SCRUB achieves the best performance, as it leverages model distillation (Pérez-Cruz, 2008), a technique well-suited for class-wise forgetting.
Appendix E Component Analysis: $\lambda_{1}$ and $\lambda_{2}$
@@
To verify the impact of each key component in the optimization problem (4) of the main paper, we analyze the nonnegative trade-off parameters $\lambda_{1}$ and $\lambda_{2}$ in fig. A3. As can be seen in fig. A3-(a) and (c), setting the retain loss regularization parameter $\lambda_{1}$ to $3$ for CIFAR-10 using ResNet-18 and $5$ for ImageNet-10 using VGG-16 in class-wise forgetting, along with the $\ell_{2}$-norm regularization parameter $\lambda_{2}=1$, enables our proposed method to achieve the highest average performance. Meanwhile, for the random data forgetting scenario, setting both $\lambda_{1}$ and $\lambda_{2}$ to $1$ yields the best average performance for CIFAR-10 using ResNet-18 and ImageNet-10 using VGG-16.
Appendix F Efficiency Analysis
MU Method Class-wise Forgetting Random Data Forgetting
CIFAR-10, ResNet-18 ImageNet-10, VGG-16 ImageNet-10, ViT CIFAR-10, ResNet-18 ImageNet-10, VGG-16 ImageNet-10, ViT
RTE Param. # RTE Param. # RTE Param. # RTE Param. # RTE Param. # RTE Param. #
Retrain $81.30$ $11.17$ $118.89$ $15.31$ $301.98$ $85.81$ $103.71$ $11.17$ $116.10$ $15.31$ $290.60$ $85.81$
FT $4.56$ $11.17$ $2.99$ $15.31$ $15.40$ $85.81$ $2.28$ $11.17$ $2.95$ $15.31$ $14.47$ $85.81$
RL $5.61$ $11.17$ $3.58$ $15.31$ $33.00$ $85.81$ $4.75$ $11.17$ $3.59$ $15.31$ $32.84$ $85.81$
GA $1.16$ $11.17$ $0.30$ $15.31$ $2.10$ $85.81$ $1.15$ $11.17$ $0.31$ $15.31$ $3.10$ $85.81$
NegGrad+ $4.72$ $11.17$ $3.21$ $15.31$ $16.70$ $85.81$ $4.22$ $11.17$ $3.17$ $15.31$ $16.40$ $85.81$
SalUn $2.81$ $5.59$ $2.56$ $7.66$ $18.73$ $42.91$ $3.01$ $5.59$ $2.85$ $7.66$ $18.19$ $42.91$
SCRUB $1.23$ $11.17$ $1.62$ $15.31$ $33.45$ $85.81$ $1.53$ $11.17$ $1.55$ $15.31$ $22.27$ $85.81$
Class-F $1.21$ $11.17$ $2.10$ $15.31$ $2.86$ $85.81$ n/a n/a n/a n/a n/a n/a
SSD $1.20$ $11.17$ $2.10$ $15.31$ $2.21$ $85.81$ $1.15$ $11.17$ $2.00$ $15.31$ $2.22$ $85.81$
Ours $2.37$ $0.03$ $6.15$ $0.15$ $4.80$ $0.15$ $6.15$ $0.03$ $6.50$ $0.15$ $8.58$ $0.15$
@@ Appendix F Efficiency Analysis
Table A2: The efficiency profile of different MU methods across two metrics for image classification under two unlearning scenarios on CIFAR-10 (RestNet-18) and ImageNet-10 (VGG-16 and ViT). Results are reported in terms of run-time efficiency (RTE) measured in minutes for the overall training phase and parameter efficiency denoted by parameter number (Param.#) in Million (M), where a smaller value is favored for each metric.
To demonstrate the efficiency profile of various MU methods under different metrics, we compare the runtime costs of forget vector calculation with those of other model-based unlearning methods, and the total number of updated parameters in the training phase. Specifically, following (Fan et al., 2023), we use run-time efficiency (RTE) as an evaluation metric, which measures the computation time of applying an MU method in minutes. To ensure a fair comparison, we report the runtime of each approximate MU method within 10 epochs since the number of iterations varies across different MU methods. Additionally, inspired by (Zhang et al., 2024b), we compare the number of trainable parameter number (Param.#). Notably, all evaluations are performed in the same computational environment with an NVIDIA A6000 GPU, ensuring fair and reliable comparisons by maintaining a consistent batch size across all MU methods. The corresponding results can be found in table A2, and we can draw the following observation: 1) By comparing the training time of model-based methods with our proposed approach using RTE within the same iterations, we observe that the forget vector method achieves competitive RTE compared to some efficient approximate model-based unlearning baselines and remains faster than retraining. Besides, it is worth noting that the forget vector method optimizes fewer parameters, owing to its input-level design and perturbation space is significantly lower-dimensional. This trade-off underscores the efficiency of our forget vector in handling parameter scaling. For instance, when comparing forgetting cases on ImageNet-10 using VGG-16 and ViT models, we find that as the parameter count increases from “15.31M” in VGG-16 to “85.81M” in ViT, the optimization time for most baseline methods increases significantly. In contrast, our method maintains a consistent efficiency level or even achieves a shorter runtime, highlighting a key advantage of our forget vector approach. 2) From a memory-efficient perspective in real-world application, existing model-based MU methods often necessitate a series of operations (e.g., fine-tuning) on the already-trained model for each forgetting request. This necessitates storing a separate model version for every request, leading to substantial storage overhead. In contrast, our approach optimizes an input-level universal perturbation, where only the “forget vector” associated with the data to be forgotten is stored with original model remains intact. Since each forget vector has the same dimensionality as the input image (e.g., 0.03M for CIFAR-10 and 0.15M for ImageNet) and is significantly smaller than the full model, our approach offers a considerable advantage in storage efficiency for practical applications.
Appendix G Additional Visualization through An Input Saliency Lens
@@
Appendix H Broader Impacts and limitations
Broader impacts.
Our study on the forget vector introduces a novel, data-driven machine unlearning approach that offers significant potential across various domains. ① Privacy Preservation and Regulatory Compliance: With increasing global regulations like GDPR mandating the right to be forgotten, our method enables effective unlearning of specific data points without requiring full model retraining, which is especially valuable for industries handling sensitive data, such as healthcare, finance, and personalized recommendation systems. By enabling compliant data removal with minimal computational overhead, our approach strengthens data privacy practices while maintaining model integrity. ② Adaptability Across Tasks: The compositional nature of forget vectors allows for the unlearning of previously unseen tasks by combining class-specific forget vectors. This adaptability extends the method’s applicability, where our method is able to respond more quickly when faced with new random unlearning tasks. ③ Efficiency in Unlearning and Model Storage: Unlike model-based MU approaches that require significant computational resources and model storage for each unlearning request, our method provides an energy-efficient alternative with less trainable parameters and low storage of optimized forget vector. This efficiency is particularly beneficial for organizations deploying large-scale AI models, where frequent updates or compliance-driven data deletions could otherwise be prohibitively expensive, contributing to the sustainability of AI development.
Limitations.
While our forget vector approach presents compelling advantages, it also comes with certain limitations that need to be addressed for broader adoption: ① Vulnerability to Adversarial Attacks: Since our method relies on input perturbations rather than direct model modifications, it may be vulnerable to white-box adversarial attacks where an attacker has access to the model and understands the forget vector mechanism. These adversaries could potentially reconstruct forgotten data or design countermeasures to bypass unlearning. Future work should focus on strengthening robustness against such attacks through adversarially hardened perturbation strategies. ② Computational Cost of Compositional Unlearning: The generation of new forget vectors through compositional operations requires access to pre-trained class-wise forget vectors. In scenarios where these vectors are unavailable, applying our method at scale could introduce overhead. ③ Lack of Theoretical Guarantees: While empirical results demonstrate the effectiveness of our method, formal theoretical guarantees on its unlearning performance and robustness remain an open challenge. Future work should focus on establishing rigorous mathematical foundations for the forget vector framework. ④ Generalization to Other Domains: Our current study focuses on image classification, and its applicability to other domains, such as natural language processing or image generation, remains to be explored. Investigating the feasibility of forget vectors in diverse machine learning tasks is an important direction for future research.